OpenStudy (anonymous):

Which expression represents in radical form?2 times x to the 5 sixths power

OpenStudy (zehanz):

If you want to know what is \[2*x^{\frac{ 5 }{ 6 } }\] in radical form, you must know what the meaning is of the fractioned exponent, 5/6. It is defined as follows:\[x ^{\frac{ a }{ b }}:=\sqrt[b]{x^a}=\left( \sqrt[b]{x} \right)^a\]So the denominator of a/b (=b) is associated with the b-th root, the nominator (a) with the a-th power. Normally, the number x is supposed to be nonnegative. This means that the order of calculation does not matter. Now try to do this for your problem.

OpenStudy (anonymous):

does that mean i have to switch it? the 5&6 @ZeHanz

OpenStudy (anonymous):

is it (2*x)the power 5/6

OpenStudy (anonymous):

yes at @gomathi

OpenStudy (zehanz):

I don't think you have to switch it, I would write it as (if gomathi is right) \[\sqrt[6]{(2x)^5}=\sqrt[6]{32x^5}\]

OpenStudy (anonymous):

32x5 so thats the answer @ZeHanz ?

OpenStudy (zehanz):

No, that would just be 32x5=160. The answer is \[\sqrt[6]{32x^5}\]

OpenStudy (anonymous):

yes 2 is also affected by the radical